### v2

# Minimum Strict Consistent Subset in Paths, Spiders, Combs and Trees

Abstract
In a connected simple graph $G=(V,E)$, each vertex of $V$ is colored by a color from the set of colors $C=\{c_{1},c_{2},\dots,c_{\alpha}\}$. We take a subset $S$ of $V$, such that for every vertex $v$ in $V\setminus S$, at least one vertex of the same color is present in its set of nearest neighbors in $S$. We refer to such a $S$ as a consistent subset (CS). The Minimum Consistent Subset (MCS) problem is the computation of a consistent subset of the minimum size. It is established that MCS is NP-complete for general graphs, including planar graphs. We expand our study to interval graphs and circle graphs in an attempt to gain a complete understanding of the computational complexity of the MCS problem across various graph classes. The strict consistent subset is a variant of consistent subset problems. We take a subset $S^{\prime}$ of $V$, such that for every vertex $v$ in $V\setminus S^{\prime}$, all the vertices in its set of nearest neighbors in $S$ have the same color as $v$. We refer to such a $S^{\prime}$ as a strict consistent subset (SCS). The Minimum Strict Consistent Subset (MSCS) problem is the computation of a consistent subset of the minimum size.
We demonstrate that MSCS is NP-hard in general graphs. We show a 2-approximation in trees. Later, we show polynomial-time algorithms in trees. Later, we demonstrate faster polynomial-time algorithms in paths, spiders, and combs.
Keywords Minimum Strict Consistent Subset  $\cdot$ MSCS for General Graphs  $\cdot$ MSCS  $\cdot$ MCS  $\cdot$ MSCS in Paths  $\cdot$ MSCS in Trees  $\cdot$ MSCS in Spiders  $\cdot$ MSCS in Combs  $\cdot$ Paths  $\cdot$ Trees  $\cdot$ Combs  $\cdot$ Spiders  $\cdot$ Nearest-Neighbor Classification  $\cdot$ Strict Consistent Subset  $\cdot$ Approximation Algorithm  $\cdot$ NP-hardness  $\cdot$ Algorithm  $\cdot$ Polynomial-time Algorithm
1 Introduction
Many supervised learning approaches use a colored training dataset $T$ in a metric space $(X,d)$ as input, with each element $t\in T$ assigned a color $c_{i}$ from the set of colors $C=\{c_{1},c_{2},\dots,c_{\alpha}\}$. The purpose is to preprocess $T$ in order to meet specific optimality requirements by immediately assigning color to any uncolored element in $X$. The closest neighbor rule, which assigns a color to each uncolored element $x$ based on the colors of its $k$ closest neighbors in the training dataset $T$ (where $k$ is a fixed integer), is a popular optimality criterion. The efficiency of such an algorithm is determined by the size of the training dataset. As a result, it is necessary to reduce the size of the training set while retaining the distance data. Hart [1] introduced this concept in $1968$ with the minimum consistent subset (MCS) problem. Given a colored training dataset $T$, the goal is to find a subset $S\subseteq T$ of the minimum cardinality such that the color of every point $t\in T$ is the same as one of its nearest neighbors in $S$. Over $2800$ citations to [1] on Google Scholar demonstrate that the MCS problem has found several uses since its inception. [2] demonstrated that the MCS problem for points in $\Re^{2}$ under the Euclidean norm is NP-complete if at least three colors color the input points. Furthermore, it is NP-complete even with two colors [3]. Recently, it was established that the MCS problem is W[1]-hard when parameterized by the output size [4].[2] and [5] offer algorithms for the MCS problem in $\Re^{2}$.
1.1 Notaions and Definitions
This paper explores the minimum consistent subset problem when $(\mathcal{X},d)$ is a graph metric. Most graph theory symbols and notations are standard and come from [6]. For every graph $G$, we refer to the set of vertices as $V$ or $V(G)$, and the set of edges as $E$ or $E(G)$. Consider a graph $G$ with a vertex coloring function $C:V(G)\rightarrow\{c_{1},c_{2},\dots,c_{\alpha}\}$. Let $C(U)$ represent the set of colors of the vertices in $U$. Formally, $C(U)=\{C(u):u\in U\}$. For any two vertices $u,v\in V(G)$, we use $\mathrm{d}(u,v)$ to signify the shortest path distance (number of edges) between $u$ and $v$ in $G$. $\mathrm{d}(v,U)=\min_{u\in U}\mathrm{d}(v,u)$ for a vertex $v\in V(G)$ and any subset of vertices $U\subseteq V$. For any graph $G$ and any vertex $v\in V(G)$, let $\mathrm{N}(v)=\{u\mbox{ }|\mbox{ }u\in V(G),(u,v)\in E(G)\}$ indicate the set of neighbors of $v$ and $\mathrm{N}[v]=\{v\}\cup\mathrm{N}(v)$. For $U\subset V$, $\mathrm{N}[U]=\cup_{u\in U}\mathrm{N}[u]$. The $r$-th neighborhood of $v$ is denoted by $\mathrm{N}^{r}[v]=\mathrm{N}[\mathrm{N}^{r-1}[v]]$, which is recursively defined. Note that $\mathrm{N}^{1}[v]=\mathrm{N}[v]$ and $\mathrm{N}[\emptyset]=\emptyset$. However, for an induced graph $G_{1}$ of $G$, the set of neighbor vertices of a vertex $v\in V(G)$ in $G_{1}$ is denoted by $\mathrm{N}_{G_{1}}(v)$ and defined by $\mathrm{N}_{G_{1}}(v)=\{u\mbox{ }|\mbox{ }u\in V(G_{1}),(u,v)\in E(G)\}$. $\mathrm{N}_{G_{1}}[v]=\{v\}\cup\mathrm{N}_{G_{1}}(v)$. Also, for $U\subset V(G_{1})$, $\mathrm{N}_{G_{1}}[U]=\cup_{u\in U}\mathrm{N}_{G_{1}}[u]$ and $\mathrm{N}_{G_{1}}^{r}[v]=\mathrm{N}_{G_{1}}[\mathrm{N}_{G_{1}}^{r-1}[v]]$. The distance between two subgraphs $G_{1}$ and $G_{2}$ in $G$ is represented as $\mathrm{d}(G_{1},G_{2})=\min\{\mathrm{d}(v_{1},v_{2})\mbox{ }|\mbox{ }v_{1}\in V(G_{1}),v_{2}\in V(G_{2})\}$. Therefore, $\mathrm{N}^{r}[v]=\{u\in V\mbox{ }|\mbox{ }\mathrm{d}(u,v)\leq r\}$. For any subset of vertices $U\subseteq V$ in a graph $G$, $G[U]$ represents the subgraph of $G$ induced on $U$. The nearest neighbor of $v$ in the set $U$ is indicated as $\mathrm{NN}(v,U)$, which is formally defined as $\{u\in U\mbox{ }|\mbox{ }\mathrm{d}(v,u)=\mathrm{d}(v,U)\}$. Also, for $X,U\subseteq V$, $\mathrm{NN}(X,U)=\cup_{v\in X}\{u\in U\mbox{ }|\mbox{ }\mathrm{d}(v,u)=\mathrm{d}(v,U)\}$. The shortest path in $G$ is $\mathrm{path}(v,u)$, which connects vertices $v$ and $u$. $\mathrm{path}(v,U)=\{\mathrm{path}(v,u)\mbox{ }|\mbox{ }u\in\mathrm{NN}(v,U)\}$.
If there is a vertex $u\in\mathrm{NN}(v,U)$ such that $C(v)=C(u)$, then $v$ is said to be covered by the set $U$. We sometimes use only $u$ covers $v$ where $u\in U$. $\mathrm{COV}(v,U)$ refers to the set of vertices in $U$ that cover $v$. $\mathrm{COV}(v,U)=\{u\in U\mbox{ }|\mbox{ }u\in\mathrm{NN}(v,U)\mbox{ and }C(u)=C(v)\}$, and for a subset $X\subseteq V$, $\mathrm{COV}(X,U)=\cup_{v\in X}\{u\in U\mbox{ }|\mbox{ }u\in\mathrm{NN}(v,U)\mbox{ and }C(u)=C(v)\}$. Thus, $u\in U$ covers $v$ indicates that $u\in\mathrm{COV}(v,U)$. If $\mathrm{COV}(v,U)=\emptyset$, then $v$ is not covered by the set $U$; otherwise, $v$ is covered by the set $U$. Similarly, if $C(X)=C(\mathrm{COV}(X,U))$, then $X$ is covered by $U$. If $v$ is not covered by the set $U$, $C(v)\neq C(\mathrm{COV}(v,U))$. For an induced graph $G(U)$, $\textup{{MCS}}(U)$ denotes the consistent subset of $G(U)$. However, simply MCS and $\textup{{MCS}}(V(G))$ are the same. Without loss of generality, we shall use $[n]$ to denote the set of integers $\{1,\ldots,n\}$. A subset $X$ is said to be a consistent subset (in short CS) for $(G,C)$ if every $v\in V(G)$, $C(v)\in C(\mathrm{NN}(v,X))$. A subset $S$ is said to be a strict consistent subset (in short SCS) for $(G,C)$ if every $v\in V(G)$, $C(v)=C(\mathrm{NN}(v,S))$. The strict consistent subset problem in graphs is defined as follows:
MCS on Graphs Input: A graph $G$, a coloring function $C:V(G)\rightarrow\{c_{1},c_{2},\dots,c_{\alpha}\}$, and an integer $\ell$. Question: Does there exist a strict consistent subset of size $\leq\ell$ for $(G,C)$?
Figure 1(A) shows an example of MCS, MSCS etc. Minimum consistent subset (MCS) refers to a consistent subset with a minimum size. For an induced subgraph $G^{{}^{\prime}}=G(U)$ with $U\subseteq V$, the minimum consistent subset of $G^{{}^{\prime}}$ is denoted by $\textup{{MCS}}(G^{{}^{\prime}})$ or $\textup{{MCS}}(U)$ and a consistent subset is denoted by $\mathcal{S}(G^{{}^{\prime}})$ or $\mathcal{S}(U)$.
Figure 1: (A) An MCS for a two-colored unweighted tree. (B) Blocks in a tree. and (C) an MCSS for the same tree.
2 Previous Results and New Results
2.1 Previous Results
Banerjee et al. [7] demonstrated that the MCS is  with when parameterized by the size of the minimum consistent set, even with only two colors. Under basic complexity-theoretic assumptions for generic networks, a FPT algorithm parameterized by $(c+\ell)$ is not feasible. This naturally raises the question of identifying the simplest graph classes for which the problem remains computationally intractable. Dey et al. [8] proposed a polynomial-time algorithm for MCS on bi-colored paths, spiders, combs, etc. Also, Dey et al. [9] proposed a polynomial-time algorithm for bi-colored trees. Arimura et al. [10] introduced a XP algorithm parameterized by the number of colors $c$, with a running time of $\mathcal{O}(2^{4c}n^{2c+3})$. Recently, we also showed that MCS on trees is NP-complete in [11]. In the article [12], we demonstrate that fixed-parameter tractability has a faster running time for the same problem. Manna presented an $(4\alpha+2)$ approximation algorithm of MCS in interval graphs and APX-hard in circle graphs in the paper [13]. A recent paper [14] introduces the concept of the minimum consistent spanning subset, which is a variant of the consistent subset problem. The restriction is that every block (a block is a connected subgraph with the same color vertices) of the graph must contain at least one vertex in the consistent subset. In the paper [14], it is assumed that a consistent subset of a graph must contain at least one vertex from each block, despite the fact that in general, a consistent subset may not contain at least one vertex from each block. They have implemented algorithms in trees that require polynomial time due to these constraints.
2.2 New Results
We demonstrate that MSCS is NP-hard in general graphs in section 3. We show a 2-approximation in trees in the same section. Later, we show $\mathcal{O}(n^{4})$ polynomial-time algorithms in trees in the section 4 for finding MSCS. In section 5, we demonstrate faster polynomial-time algorithms in paths, spiders, and combs with running times $\mathcal{O}(n)$, $\mathcal{O}(n^{2})$, $\mathcal{O}(n^{3})$, respectively for finding MSCS.
3 Preliminaries, NP-hardness in General Graphs and 2-Approximation Algorithm in Trees
3.1 Preliminaries
Observation 3.0.1.
If all vertices of $T$ have the same color, i.e. $T$ is monochromatic, then every vertex of $T$ is a minimum strict consistent subset for $T$.
The definition of block as a maximal subset of connected vertices of the same color in $\mathrm{T}$. In view of Observation 3.0.1 we may assume that $\mathrm{T}$ has more than one block. Let $k\geq 2$ be the number of blocks in $\mathrm{T}$. We define the block tree of $\mathrm{T}$, denoted by $\mathcal{B}(\mathrm{T})$, as the tree with $k$ vertices, each corresponding to a block of $\mathrm{T}$, and there is an edge between any two vertices if their corresponding blocks are neighbors in $\mathrm{T}$. In other words, $\mathcal{B}(\mathrm{T})$ is obtained by contracting all blocks of $\mathrm{T}$. Notice that each vertex of $\mathcal{B}(\mathrm{T})$ corresponds to a block of $\mathrm{T}$, and vice versa. We refer to a block of $\mathrm{T}$ as a leaf block if its corresponding vertex in $\mathcal{B}(\mathrm{T})$ is a leaf. A block of $\mathrm{T}$ that is not a leaf block is called an internal block (see Figure 2(A) and 2(B)).
Figure 2: (A) Example of a tree $\mathrm{T}$ together with its blocks. (b) The block tree $\mathcal{B}(\mathrm{T})$ of $\mathrm{T}$.
Lemma 3.1.
Every strict consistent subset has at least one vertex from each block of $T$.
Proof.
Let $B_{i}$ be the block such that no vertex of $B_{i}$ appear in a strict consistent subset. As no vertex of $B_{i}$ is a strict selective vertex, then it is obvious that all the vertices of $B_{i}$ must be strictly covered by some of the vertices of a different block $B_{j}$ ($i\neq j$ and $C(B_{i})=C(B_{j})$). Let $v$ be a vertex in $B_{i}$, which is strictly covered by a vertex $u$ of $S_{j}$. We must have the shortest path (say $P$) between $v$ and $u$ such that $P$ must have at least vertex $w$ with a different colour than the colour of $S_{i}$; otherwise, $S_{i}$ and $S_{j}$ are identical, and our lemma is proved. As $C(w)\neq C(S_{i})$, then $u$ is one of the nearest neighbors of $w$ and $w$ is not strictly covered because $C(u)\neq C(w)$. Now, if $w$ is contained in the strict consistent subset, then $v$ is not again strictly covered by $u$ because $dist(v,w)<dist(v,u)$ and $c(v)\neq c(w)$; a contraction.
Lemma 3.2.
Any minimum strict consistent subset of a tree $T$ contains exactly one vertex from each leaf block of $T$.
Proof.
We use contradiction to prove this lemma. Consider a strict minimum consistent subset $S^{\prime}$ that contains more than one vertex from some leaf block, say $B$. Let $\mathrm{N}$ be the neighboring block of $B$. Let $x$ be a vertex of $S^{\prime}$ in $B$ that is closest to $\mathrm{N}$, as in Figure 3. Let $S^{\prime\prime}$ be the set obtained from $S^{\prime}$ by removing all the vertices of B except for $x$. We claim that $S^{\prime}$ is a strict consistent subset for $\mathrm{T}$; this would contradict $S^{\prime}$ being minimum.
To prove the above claim, it suffices to show that $x$ is a vertex of $S^{\prime}$ that is closest to every vertex of $B$. Let $t$ be a vertex of $S^{\prime}$ in $\mathrm{N}$ that is closest to $B$. Let $v$ be the last vertex of $B$ on the path from $x$ to $t$, as depicted in the figure to the right. As $S^{\prime}$ is a strict consistent subset, we have $\mathrm{d}(v,x)<\mathrm{d}(v,t)$. Now consider any vertex in $u\in B$. The path between $u$ and $t$ passes through $v$. Therefore $\mathrm{d}(u,x)<\mathrm{d}(u,t)$, which proves the claim. ∎
Figure 3: Illustration of Lemma 3.2.
3.2 NP-hardness of the MSCS in General Graphs
Our reduction is from dominating set problem on a graph $G=(V,E)$. From $G$, we construct another graph $G^{\prime}=G_{1}\cup G_{2}\cup\{v_{spc}\}$ where $G_{1}$ and $G_{2}$ are the copies of $G$ and $v_{spc}$ is a singleton node. Each edge of $G_{1}$ and $G_{2}$ has weight $1$. Every vertices of $G_{1}$ are directly connected by edges with all the vertices of $G_{2}$ with weights of edge $3-4\epsilon$ where $0<\epsilon<\frac{1}{2}$. Additionally, all vertices of $G_{2}$ are directly connected by edges of weight $2\epsilon$ to $v_{spc}$. All the vertices of $G_{1}$ are colored red and all the vertices of $G_{2}$ are colored blue and $v_{spc}$ is also colored blue. This completes the construction.
Lemma 3.3.
$G$ has a dominating set of size $k$ if and only if $G^{\prime}$ has a strict consistent subset of size $k+1$.
Proof.
Suppose that $G$ has a dominating set $S$ of size $k$. We will show that $G^{\prime}$ has a strict consistent subset of size $k+1$. Suppose that $S^{\prime}\leftarrow\emptyset$. For each vertex of $S$, we put the corresponding vertex in $G_{1}$ in $S^{\prime}$. Also, we put $v_{spc}$ in $S^{\prime}$. Therefore, $\lvert S^{\prime}\rvert=k+1$. We claim that $S^{\prime}$ is a strict consistent subset of $G^{\prime}$. For a red color vertex $v\in G_{1}$, if $v\notin S^{\prime}$, then $\mathrm{d}(v,v_{spc})=3-4\epsilon+2\epsilon>1$, but $v$ has a neighbor vertex in $G_{1}$ which is in $S^{\prime}$ of length $1$. Therefore, all the red-colored vertices are strictly covered. Also, for a blue-colored vertex $u\in G_{2}$, $\mathrm{d}(u,G_{1})=3-4\epsilon>2\epsilon=\mathrm{d}(u,v_{spc})$. Therefore, all the red-colored vertices are strictly covered by $v_{spc}$.
Again, suppose that $S^{\prime}$ is a strict consistent subset of $G^{\prime}$ such that $\lvert S^{\prime}\rvert=k+1$. Therefore $v_{spc}$ must included in $S^{\prime}$; otherwise, size of $S^{\prime}$ must be greater than $k+1$. Therefore, all the vertices of $G_{2}$ are strictly covered. Hence we have $k$ vertices from $G_{1}$ in $S^{\prime}$. We take corresponding vertices of $G_{1}$ in $S^{\prime}$ into a set $S$. We claim that $S$ is the dominating set of $G$. If not then there exists a vertex $v\in S$ such that there is no neighbor of $v$ in $S$ and $v\notin S$. Therefore the corresponding vertex of $v$ in $G_{1}$ can not be covered strictly by $S^{\prime}$; a contradiction. ∎
3.3 2-Approximation Algorithm of MSCS in Trees
Recall the block tree $\mathcal{B}(\mathrm{T})$; see Figure 2. Let $S$ be the empty set. For every edge $e\in\mathcal{B}(\mathrm{T})$ our algorithm adds to $S$ the two vertices of $\mathrm{T}$ that correspond to the endpoints of $e$. It is easily seen that $S$ is a feasible solution for the MSCS problem. We claim that $S$ is a 2-approximate solution. Let $b$ denote the number of blocks in $\mathrm{T}$, which is the same as the number of vertices of $\mathcal{B}(\mathrm{T})$. Since $\mathcal{B}(\mathrm{T})$ is a tree, it has $b-1$ edges. Therefore $\lvert S\rvert=2b-2$. On the other hand the size of any optimal solution $OPT$ is at least $b$ because it must contain at least one vertex from each block of $\mathrm{T}$ (by Lemma 3.1), and hence $\lvert OPT\rvert\geq b$. Therefore
$\lvert S\rvert=2b-2\leq 2\lvert OPT\rvert-2.$
We note that this elementary analysis is the best possible for this algorithm, for example if $\mathrm{T}$ is a path with blocks of size three, then our algorithm picks two vertices for each block (except for the two leaf blocks) while the optimal solution picks one vertex (the middle vertex) from each block.
4 Algorithm for Finding MSCS in Trees
The Lemma 3.1 clearly demonstrates that at least one vertex from each block must appear in any strict consistent subset which is the restriction of the paper [14]. However, the definitions of minimum consistent spanning subset and minimum strict consistent subset are something distinct, but the Lemma [14] is applicable for these two variants of MCS. Therefore, we use a very similar argument from the paper [14] as follows for our MSCS problem.
Lemma 3.2 suggests a more constrained version of the MSCS problem, in the sense that we can fix a leaf block B and enforce exactly one vertex of B to be in the solution. As we do not know in advance which vertex of B is in the optimal solution, we try all of them and report the best answer.
Our algorithm employs a nontrivial dynamic programming approach. First we introduce the subproblems that will be generated throughout the algorithm and then we will show how to solve the subproblems recursively.
4.1 Defining Subproblems For Solving MSCS in Trees
We denote each subproblem by $\mathrm{T}(x,y)$ where $x$ and $y$ are two given vertices of $\mathrm{T}$. Consider $\mathrm{path}(x,y)$ between $x$ and $y$ in $\mathrm{T}$ and let $u\in\mathrm{N}(y)$ in $\mathrm{path}(x,y)$ (it might be the case that $x$ = $u$). By removing the edge $(u,y)$ from $\mathrm{T}$ we obtain two subtrees. Let $\mathrm{T}_{y}$ be the subtree containing $y$ (see Figure 4(A)). Let $\mathrm{T}^{\prime}$ be the union of $\mathrm{path}(x,y)$ and $\mathrm{T}_{y}$ (see Figure 4(B)). We define $\mathrm{T}(x,y)$ to be the MSCS problem on $\mathrm{T}^{\prime}$ with the following constraints:
  • $x$ must be in the solution, and
  • all the vertices from $x$ to $u$ on $\mathrm{path}(x,y)$ must be covered by $x$.
These constraints imply that the vertices from $x$ to $u$ should have the same color.
Figure 4: (A) The tree $\mathrm{T}$ , and (B) the tree $\mathrm{T}^{\prime}$.
4.2 Solving Subproblems For Solving MSCS in Trees
We denote the size of the (constrained) MSCS for $\mathrm{T}(x,y)$ by $S(x,y)$. If $\mathrm{T}(x,y)$ has no solution, then we set $S(x,y)=+\infty$. To solve $\mathrm{T}(x,y)$ we proceed as follows
If $\mathrm{T}^{\prime}$ is monochromatic, then (by Observation 3.0.1) we return $x$ as the solution. In this case, $S(x,y)=1$. Assume that $\mathrm{T}^{\prime}$ is not monochromatic. We root $\mathrm{T}^{\prime}$ at $x$.
Lemma 4.1.
If $\mathrm{T}(x,y)$ has a solution, then any solution of $\mathrm{T}(x,y)$ contains a vertex $z$ in the same block as $x$ or in a neighboring block of $x$ such that all vertices on $\mathrm{path}(x,z)$ are covered by $x$ or by $z$.
Proof.
As $\mathrm{T}^{\prime}$ is multicolored, any solution of $\mathrm{T}(x,y)$ should contain at least two vertices. In particular it should contain at least one vertex from each neighboring block of $x$. In any solution of $\mathrm{T}(x,y)$ a vertex that is closest to $x$ satisfies the statement of the lemma for $z$. ∎
Let $z$ be any vertex of $\mathrm{T}^{\prime}$ that satisfies the constraints of Lemma 4.1 (see Figure 5) (It might be the case that $z=y$. Also $z$ could be in $x$’s block or in $x$’s neighboring block.) If such a vertex $z$ does not exist then $\mathrm{T}(x,y)$ has no solution and thus we set $S(x,y)=+\infty$.
Figure 5: Solving $\mathrm{T}(x,y)$ recursively in terms of $\mathrm{T}(x,v^{\prime})$ and $\mathrm{T}(z,v^{\prime})$ where $z$ is a valid pair for $x$.
By Lemma 4.1 all vertices of $\mathrm{path}(x,z)$ are covered by $x$ or $z$. Since $u$ must be covered by $x$ (as imposed by the definition of $\mathrm{T}(x,y))$, we must have $\mathrm{d}(u,x)<\mathrm{d}(u,z)$, and thus $\mathrm{d}(z,x)>2\cdot\mathrm{d}(u,x)$. Moreover as described before $\mathrm{path}(x,z)$ has $2k$ vertices (including a and z). Therefore, the first $k$ vertices must be covered by $x$ and the second $k$ vertices by $z$.
Now we are in a problem setting where both $x$ and $z$ must be in the solution, and all vertices on $\mathrm{path}(x,z)$ must be covered by $x$ and $z$. We denote this more constrained version of $\mathrm{T}(x,y)$ by problem $\mathrm{T}(x,y,z)$ (we do not call this a subproblem for a reason to be determined later). In other words $\mathrm{T}(x,y,z)$ is the MSCS problem on $\mathrm{T}^{\prime}$ with the following constraints:
  • $z$ is in the same block as $x$ or in a neighboring block of $x$,
  • $x$ and $z$ must be in the solution,
  • $\mathrm{d}(z,x)>2\cdot\mathrm{d}(u,x)$, and
  • $\mathrm{path}(x,z)$ has $2k$ where the first k vertices must have the same color as $x$ and the second $k$ vertices must have the same color as $z$.
We refer to any vertex $z$ that satisfies the above constraints as a valid pair for $x$. Now we show how to solve $\mathrm{T}(x,y,z)$. We denote the size of the solution for $\mathrm{T}(x,y,z)$ by $S(x,y,z)$. Let $X$ be the set of all the vertices on $\mathrm{path}(y,z)$ that are closer to $x$ than to $z$ as in Figure 5. Let $Z$ be the set of all the vertices on $\mathrm{path}(y,z)$ that are closer to $z$ than to $x$.
To solve $\mathrm{T}(x,y,z)$ we define two sets $\mathcal{X}$ and $\mathcal{Z}$ as follows. For each vertex $v$ in $X$, we add to $\mathcal{X}$ all children of $v$ that are not on the path $\mathrm{path}(y,z)$. For each vertex $v$ in $Z$, we add to $\mathcal{Z}$ all children of $v$ that are not on the path $\mathrm{path}(y,z)$. Then, the solution of $\mathrm{T}(x,y,z)$ is obtained by taking the union of $\{x,z\}$ with the solutions of $\mathrm{T}(x,v^{\prime})$ for all $v^{\prime}\in\mathcal{X}$ and the solutions of $\mathrm{T}(z,v^{\prime})$ for all $v^{\prime}\in\mathcal{Z}$. It might happen that some choices of $z$ does not lead to a valid solution, but then a subsequent subproblem, e.g. $\mathrm{T}(z,v^{\prime})$ will be infeasible and $S(z,v^{\prime})=+\infty$.
One might wonder that for two vertices $v^{\prime}_{1},v^{\prime}_{2}\in\mathcal{Z}$ the solutions of $\mathrm{T}(z,v^{\prime}_{1})$ and $\mathrm{T}(z,v^{\prime}_{2})$ may affect each other. However, this cannot happen because by the definition of $\mathrm{T}(.,.)$ all vertices on paths from $z$ to the parents of $v^{\prime}_{1}$ and $v^{\prime}_{2}$ must be covered by $z$, and thus any vertex in the solutions of $\mathrm{T}(z,v^{\prime}_{1})$ and $\mathrm{T}(z,v^{\prime}_{2})$ lies in the same level as $z$ or in a lower level (i.e. higher depth) in $\mathrm{T}^{\prime}$. The same argument applies to vertices in $\mathcal{X}$. Notice that no vertex has an equal distance to multiple vertices of different colors in the final solution because of the definition of MSCS. Therefore,
$S(x,y,z)=2+\sum_{v^{\prime}\in\mathcal{X}}S(x,v^{\prime})+\sum_{v^{\prime}\in\mathcal{Z}}S(z,v^{\prime})-\lvert\mathcal{X}\rvert-\lvert\mathcal{Z}\rvert,$
where the subtractive terms $\lvert\mathcal{X}\rvert$ and $\lvert\mathcal{Z}\rvert$ come from the fact that $x$ and $z$ are counted in each $S(x,v^{\prime})$ and $S(z,v^{\prime})$. Thus, we are able to solve $\mathrm{T}(x,y,z)$ in terms of $\mathrm{T}(x,v^{\prime})$ and $\mathrm{T}(z,v^{\prime})$ which are smaller instances of $\mathrm{T}(x,y)$.
Now we turn our attention back to the original problem $\mathrm{T}(x,y)$. If we knew $z$ then the solution of $T(x,y,z)$ would also be a solution for $T(x,y)$. But we do not know $z$ in advance. Therefore we consider all valid pairs $z$ for $x$ and take one that gives the smallest solution. Let $P$ denote the set of all valid pairs for $x$. Then
$S(x,y)=\min\{S(x,y,z)\text{ }|\text{ }z\in P\}.$
If $P$ is the empty set, i.e. $x$ has no valid pairs, then we set $S(x,y)=+\infty$. This is the end of our solution for the subproblem $\mathrm{T}(x,y)$; the problem $T(x,y,z)$ was introduced to simplify the description of our recursive solution.
4.3 Solving the Original Problem in Trees
In this section we show how to solve the original MSCS problem on the tree $\mathrm{T}$. Let $S$ denotes an optimal solution for this problem. If $\mathrm{T}$ is monochromatic then we return an arbitrary vertex as a solution. Assume that $\mathrm{T}$ has more than one colors. Let $L$ be a leaf block of $tree$, and let $\mathrm{N}$ be the (only) neighboring block of $L$. Let $y$ be the vertex of $\mathrm{N}$ adjacent to a vertex of $L$. By Lemma 3.2, any optimal solution has exactly one vertex, say $x$, from L. Thus, $x$ is the closest vertex of $S$ to all vertices in $L$. Hence the original problem is an instance of the $\mathrm{T}(x,y)$ problem that was introduced earlier. Since we do not know $x$, we try all vertices of $L$. Therefore,
$\lvert S\rvert=\min\{S(x,y)\text{ }|\text{ }x\in L\}.$
The correctness of our algorithm is implied by Lemma 3.2, the fact that in any optimal solution, all the vertices of $L$ are closer to $x$ than to any other vertex of $S$, and from the correctness of our solution for $\mathrm{T}(x,y)$.
4.4 Running Time Analysis
In this section we analyze the running time of our algorithm for the MSCS problem on a tree $\mathrm{T}$ with $n$ vertices. The algorithm follows a top-down dynamic programming approach and consists of subproblems of the form $\mathrm{T}(x,y)$ where $x$ and $y$ are two vertices of $\mathrm{T}$. In total we have $\mathcal{O}(n^{2})$ subproblems of this form. To solve each subproblem we examine $\mathcal{O}(n)$ valid pairs for $x$. For each valid pair $z$ of $x$ we look up to solutions of $\mathcal{O}(n)$ smaller subproblems, namely $\mathrm{T}(x,v^{\prime})$ and $\mathrm{T}(z,v^{\prime})$. Thus each subproblem $\mathrm{T}(x,y)$ can be solved in $\mathcal{O}(n^{2})$ time. Therefore the total running time of the algorithm is $\mathcal{O}(n^{4})$.
One might wonder that for any fixed pair $(x,y)$ if the size of solutions of $T(x,y,z)$, i.e. $S(x,y,z)$, is monotonic (increasing or decreasing) with respect to $\mathrm{d}(x,z)$ then we could save an $\mathcal{O}(n)$ factor in the computation of $\mathrm{T}(x,y)$. However, this is not always the case; Figure 6 shows that if we increase $\mathrm{d}(x,z)$ then $S(x,y,z)$ could increase (Figure 6(A)) or decrease (Figure 6(B)).
Figure 6: If we increase $\mathrm{d}(x,z)$ then $S(x,y)$ can increase (A) or decreases (B)
4.5 Extension to Weighted Trees
Our algorithm can simply be extended to solve the MSCS problem on weighted trees within the same time complexity. The only parts that need to be adjusted are the definitions of distances and valid pairs. The distance $\mathrm{d}(x,y)$ between two vertices $x$ and $y$ is now the the shortest distance in the weighted tree. The existence of a valid pair $z$ for a (Lemma 3.2) can be shown by taking a vertex in the solution whose weighted distance to $x$ is minimum. The validity of $z$ is now verified by similar constraints, except for the last constraint, which we adjust as follows:
  1. 1.
    All vertices on path $\mathrm{path}(x,z)$ that are closer to $x$ (in terms of weighted distance) must have the same color as $x$ and all vertices that are closer to $z$ must have the same color as $z$.
5 Algorithm for Finding MSCS in Paths, Spiders, Combs
5.1 Algorithm for Finding MSCS in Paths
We employ arguments that are quite similar to those of the paper [8], which addressed the algorithm of MCS in paths. As a result, the algorithm for MSCS in paths is as follows.
In a path graph $G$, the vertices in $V(G)$ are listed in the order $p_{1},p_{2},\dots,p_{n}$. Each pair of consecutive vertices define an edge of the graph, i.e., $E(G)=\{(p_{i},p_{i+1}),i=1,2,\dots,n-1\}$. Each vertex has degree two except the two terminal vertices of the path that have degree one. In a $\alpha$-chromatic path graph, each vertex is assigned with one color in $\{c_{1},\dots,c_{\alpha}\}$. The blocks of the path are shown in Figure 7. In a block of length 1, the vertex itself is always selected.
Figure 7: Blocks of a path
Lemma 5.1.
In the minimum strict consistent subset of a path graph, each block will have at least one and at most two vertices in $S$. Moreover, exactly one vertex will be sufficient from the first and the last block.
Proof.
The presence of at least one element from each block in $S$ is obvious by using Lemma 3.1. It is also obvious that if more than two vertices from a block $B$ are chosen in $S$ then all the chosen vertices in that block excepting the first and last one can be dropped without violating the strict consistency, and thereby reducing the size of $S$. By the same argument if more than one vertex from the first (resp. last) block is chosen in $S$ then all those chosen vertices excepting the rightmost in the first block (resp. leftmost in the last block) can also be dropped without violating the consistency. ∎
Algorithm: Consider a pair of adjacent blocks $B_{j}$ and $B_{j+1}$. Assume, without loss of generality, $\lvert B_{j}\rvert\leq\lvert B_{j+1}\rvert$. For each member $p_{i}\in B_{j}$, there exists at most one vertex, say $p_{k}\in B_{j+1}$, such that if $p_{i}$ is included in $S$ then $p_{k}$ of $B_{j+1}$ must be included in $S$ to satisfy the strict consistency property of the boundary vertices of $B_{j}$ and $B_{j+1}$ that are adjacent to each other (see Figure 8). Thus, $(pi,p_{k})$ forms a valid-pair. Now, we define the overlay graph $H$ as follows. The vertices of $H$ are the vertices of $G$, and $k$ dummy vertices $D=\{d_{1},\dots,d_{k}\}$, where $k$ is the number of blocks. Therefore, $V(H)=V(G)\cup D$. The edges in the set $E(H)$ are of two types. For each valid-pair, we add a directed type-1 edge in $V(H)$. For each vertex $p_{i}$ in a block $B_{j}$ we add two directed type-2 edges $(p_{i},d_{j})$ and $(d_{j},p_{i})$ in $V(H)$ The weight of each type-1 edge is 00. The type-2 edges incident to $D\backslash\{d_{1},d_{k}\}$ have weight $1$. Each type-2 edge incident to ${d_{1},d_{k}}$ has weight 00. For the complete demonstration of the graph $H$. A forward $s-t$ path is a path from $s$ to $t$ where the indices of the $p_{i}$ vertices appear in increasing order. Now we find the shortest forward $s-t$ path with $s=d_{1}$ and $t=d_{k}$ in the graph $H$, and remove the $d_{j}$’s to obtain MSCS of the original path graph $G$.
Figure 8: Valid pair $(p_{i},p_{k})$
Theorem 5.2.
The shortest $s-t$ path of the overlay graph $H$ gives the minimum strict consistent subset of the path G, and it executes in $\mathcal{O}(n)$ time.
Proof.
We first prove that any forward $s-t$ path of the graph $H$ constructed from the given path graph $G$ gives a strict consistent subset of the graph $G$. Observe that at least one vertex from each block is present in the $s-t$ path of the graph H. The reason is that the edges corresponding to the valid pairs are defined only between adjacent blocks in forward direction, and each dummy vertex (say $d_{i}$) is bidirectionally connected with only the vertices of one block in the graph $H$. Each edge $(v_{i},v_{j})$ of $H$ between two consecutive blocks in $G$ justify the strict consistency of the vertices $\{v_{i},v_{i+1},\dots,v_{j}\}$ of the path graph $G$. As all the vertices in any path of $H$ between a pair of vertices $v_{i}$, $v_{j}$ in the same block are of same color, choice of those vertices on the path in the strict consistent subset does not destroy the consistency property of the not chosen vertices between $v_{i}$ and $v_{j}$. Thus, any $s-t$ path with $s=d_{1}$ and $t=d_{r}$ in $H$ gives a strict consistent subset of the vertices in the path graph $G$.
Now, we will consider the nature of the minimum $s-t$ path in $H$. Each forward move in the minimum $s-t$ path from a vertex $v_{i}\in B_{j}$ either reaches a vertex $v_{\ell}\in B_{j+1}$ or to a vertex $v_{m}\in R_{j}$ through the dummy vertex $d_{j}$ where $i<m$. If no dummy vertex in $\{d_{2},\dots,d_{k-1}\}$ is visited in the $s-t$ path then exactly one vertex is present from each block in the obtained path. However, the presence of every dummy vertex $d_{i}$ in the shortest $s-t$ path implies that two vertices of the corresponding block is present in the consistent subset obtained by that $s-t$ path.
Also, the construction of the graph suggests that if a subset of vertices in $V(H)\cup D$ do not form a path, then they cannot form a consistent subset. The minimality in the size of the consistent subset is justified from the choice of the shortest $s-t$ path.
The number of type-1 edges in the graph $H$ is at most $n$ since each vertex in a block $G$ can participate in at most $1$ valid pair in its succeeding block. The number of type-2 edges is $2n$ since each vertex in $G$ is bidirectionally connected with the dummy vertex of its corresponding block. Thus we have $\mathcal{O}(n)$ edges in $H$. In the special case of integer weights and directed connected graphs, Dijkstra’s algorithm for shortest path executes in $\mathcal{O}(|E(G)|)$ time [15]. Thus, the time complexity follows. ∎
Remark: A minor tailoring of the same algorithm works for a $\alpha$-chromatic cycle graph $G$, where the vertices in $V(G)$ are connected in a closed chain.
5.2 Algorithm for Finding MSCS in Spiders
Let $B$ denote the block containing the center of $\mathrm{T}$. Any solution must contain some vertex from $B$, among which one is closest to the center. We examine each vertex $b\in B$ as a vertex in the solution that is closest to the center. Then we break the tree into some paths each of which is defined by two endpoints that are $b$ and a leaf of $\mathrm{T}$. We solve the MSCS problem on each path independently (with the constraint that $b$ must be in the solution) and then take the union of the solutions as a solution for the MSCS problem on $T$. The total running time is $\lvert B\rvert\cdot\mathcal{O}(n)=\mathcal{O}(n^{2})$.
5.3 Algorithm for Finding MSCS in Combs
We use our algorithm of previous section (for general trees) and show how to save a linear factor from the running time.
Recall from subsection 4.1 that we solve $\mathrm{T}(x,y,z)$ in linear time by looking up to solutions of $\mathcal{O}(n)$ smaller subproblems of the form $\mathrm{T}(x,v^{\prime})$ and $\mathrm{T}(z,v^{\prime})$. We are going to solve this in constant time with the help of pre-processing. We take advantage of the fact that each vertex $v$ on the path $\mathrm{path}(x,z)$ has at most one child that can be $v^{\prime}$ (because $\mathrm{T}$ is a comb) and the fact that each subproblem $T(x,v^{\prime})$ is an MSCS instance on a path (again because $\mathrm{T}$ is a comb).
Let $s_{1},\dots,s_{k}$ be the vertices of the skeleton in this order. For each $i\in\{1,\dots,k\}$ let $l_{i}$ be the leaf of $\mathrm{T}$ in the dangling path at $s_{i}$. We define two $n\times k$ matrices $\mathcal{P}$ and $\mathcal{Q}$ where the rows represent the $n$ verties of $\mathrm{T}$ and the columns represent the $k$ vertices of the skeleton. For each vertex $x\in\mathrm{T}$ and each $i\in\{1,\dots,k\}$ let $l_{i}$ the entry $\mathcal{P}[x][i]$ stores the size of the solution for the MSCS problem on the path from $x$ to $l_{i}$ with the constraint that $a$ is in the solution and all vertices from $a$ to $s_{i}$ are covered by $x$ and this can be computed in linear time using path algorithm. Each entry $\mathcal{Q}[x][i]$ stores the sum of $\mathcal{P}[a][j]$ for all indices $j$ of skeleton vertices that lie on the path $\mathrm{path}(x,s_{i})$. Thus $\mathcal{P}$ and $\mathcal{Q}$ can be computed in $\mathcal{O}(kn^{2})$ and $\mathcal{O}(k^{2}n)$ time, respectively, in a pre-processing step.
Let $m$ be the index such that all vertices on the path $\mathrm{path}(x,p_{m})$ have the same color as $x$ and all vertices on the path $\mathrm{path}(p_{m+1},z)$ have the same color as $z$ (this corresponds to definitions of the sets $A$ and $Z$ in the original algorithm). Then $\mathcal{Q}[x][m]=\sum_{v^{\prime}\in\mathcal{A}}S(x,v^{\prime})$ and $\mathcal{Q}[z][m+1]=\sum_{v^{\prime}\in\mathcal{Z}}S(z,v^{\prime})$. Since $\mathcal{Q}[x][m]$ and $\mathcal{Q}[z][m+1]$ are already computed, we can solve $\mathrm{T}(x,y,z)$ in constant time. Therefore, the total running time of the algorithm is $\mathcal{O}(n^{3})$.
6 Concluding Remarks
As the paper MSCS is a variant of the paper MCS, therefore, the algorithms for MSCS problems in other graph classes are open. Moreover, as MSCS is NP-hard, so the approximation algorithm, FPT, APX-hard, PTAS are also open for general graphs.